# Positive microlocal holonomies are globally regular

### v1

@@ 1. Introduction
Specifically, our main result is to show that such regular extensions exist if the relative cycle intersects positively an $\mathbb{L}$-compressing system. It applies to arbitrary Legendrian links and Lagrangian fillings (and arbitrary microlocal rank), both generalizing and independently recovering our previous results on global regularity of cluster variables. A key part of our argument is the construction of Hochschild cycles for the aforementioned category that lift these microlocal merodromies. The proof highlights the advantages of working within the dg-categorical framework when studying Legendrian links. Even in known cases, our argument bypasses previously required algebraic computations needed to argue regularity and, using the derived stack of pseudo-perfect objects, provides a clearer categorical argument for regularity. In brief, apart from its greater level of generality, a strength of the result is that it provides new conceptual and geometric reasons for the global regularity of such functions.
@@ 1.1. Scientific context
Let $\Lambda\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{\text{st}})$ be a Legendrian link in the ideal contact boundary of the cotangent bundle of $\mathbb{R}^{2}$, e.g. any Legendrian link in a contact Darboux chart. The study of Lagrangian fillings of $\Lambda$ is a pillar of low-dimensional contact topology, see e.g. [BST15, EENS13, EHK16, EP96, EN22, GKS12, Lev16, NRS+20, Pol91, STZ17]. Recently, the results from [CG22, CG24] and the trilogy [CGG+22, CW24, CZ22] provided a better understanding of the moduli space $\mathfrak{M}(\Lambda)$ of Lagrangian fillings and its geometric structures. Note that, in addition to new results in contact topology, studying these moduli spaces $\mathfrak{M}(\Lambda)$ from this contact topological perspective has lead to a number of applications to cluster algebras, see e.g. [CGGS20, CGG+22, CLSBW23, CW24]. Now, a key technique to understand such moduli spaces $\mathfrak{M}(\Lambda)$ is the construction of regular functions on them. The main result of this article is a general construction of such regular functions coming from a microlocal version of parallel transport.
In a nutshell, from [CW24, Section 4], certain relative 1-cycles $\eta\in H_{1}(L,\Lambda)$ in a Lagrangian filling $L$ (related to $\mathbb{L}$-compressing systems) allow us to define regular functions $A_{\eta}:T(L)\longrightarrow\mathbb{C}$ on an open subset $T(L)\subset\mathfrak{M}(\Lambda)$. If $\Lambda$ is particular enough and certain combinatorics are present, e.g. weaves in the case of [CGG+22], grid plabic graphs in the case of [CW24] and 3D plabic graphs in the case of [GLSBS22], one can sometimes argue that $A_{\eta}$ extends to a regular function on the entirety of $\mathfrak{M}(\Lambda)$. This has significant consequences, including the construction of cluster structures [CGG+22, CW24], holomorphic symplectic structures [CGGS20], the existence of Lagrangian fillings in each cluster seed [CG24] and geometric realizations of Donaldson-Thomas transformations [CLSBW23, CW24].
@@ 1.2. Main result
Let $\Lambda\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{st})$ be a Legendrian link and $\mathfrak{t}\subset\Lambda$ a set of basepoints, with at least one basepoint per component of $\Lambda$. We say $\Lambda$ is $\mathfrak{t}$-pointed if such a choice $\mathfrak{t}$ of basepoints has been made. In Section 2 we introduce a refinement $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ of the dg-category $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})$ of sheaves in $\mathbb{R}^{2}$ with singular support on $\Lambda$ which takes into account the microstalks at $\mathfrak{t}$. By construction, $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ is a smooth dg-category and we denote by $\mathfrak{M}({\Lambda,\mathfrak{t}})$ its moduli of pseudo-perfect objects, cf. [TV07, Section 3.1]. The derived stack $\mathfrak{M}({\Lambda,\mathfrak{t}})$ is the model we use for the moduli space of Lagrangian fillings of $(\Lambda,\mathfrak{t})$: intuitively, its classical closed points parameterize Lagrangian fillings of $\Lambda$ endowed with local systems. For precise details, see e.g. [CL23, CW24] and [TV07]. The construction of [TV07] is such that there exists a map
@@
Let $T\subset\Lambda\setminus\mathfrak{t}$ be another set of basepoints such that there exists exactly one basepoint of $\mathfrak{t}$ in each interval of $\Lambda\setminus T$. We say $\Lambda$ is $(\mathfrak{t},T)$-pointed if such choices of basepoints $\mathfrak{t}$ and $T$ have been made. Consider an embedded exact Lagrangian filling $L\subset(T^{*}\mathbb{R}^{2},\lambda_{\text{st}})$ of $\Lambda$ endowed with an $\mathbb{L}$-compressing system $\mathscr{D}=\{\gamma_{1},\ldots,\gamma_{b_{1}(L)}\}$, as defined in [CG24, Section 4]. Succinctly, a (complete) $\mathbb{L}$-compressing system is a set of disjoint Lagrangian disks $\{D_{1},\ldots,D_{b_{1}(L)}\}$, each properly embedded in the complement $T^{*}\mathbb{R}^{2}\setminus L$ of $L$ with immersed boundaries $\gamma_{i}:=\partial\overline{D_{i}}\subset L$ and such that the union $\mathbb{L}:=L\cup(D_{1}\cup\ldots D_{b_{1}(L)})\subset T^{*}\mathbb{R}^{2}$ is a relative Lagrangian skeleton for $(T^{*}\mathbb{R}^{2},\Lambda)$. See also [CW24, Section 1]. [^1] The assumption that $\mathbb{L}:=L\cup(D_{1}\cup\ldots D_{b_{1}(L)})\subset T^{*}\mathbb{R}^{2}$ is a relative Lagrangian skeleton for $(T^{*}\mathbb{R}^{2},\Lambda)$ is not explicitly stated in [CW24, Section 1], but all known examples of $\mathbb{L}$-compressing systems, in the sense of [CW24], satisfy this assumption. By definition, a relative cycle $\eta:[0,1]\longrightarrow(L,\mathfrak{t})$ is said to be $\mathscr{D}$-positive if $\langle\eta,\gamma_{i}\rangle\geq 0$ for all $i\in[1,b_{1}(L)]$.
Let ${T}(L)\subset\mathfrak{M}({\Lambda,\mathfrak{t}})$ be the open set ${T}(L)\cong\mathrm{Loc}(L,T)$ given by the embedded exact Lagrangian filling $L$, see e.g. [CL23, Appendix B] or [CW24, Section 2] and references therein. By construction, the microlocal monodromy $A_{\eta}\in\Gamma({T}(L),\mathcal{O}_{{T}(L)})$ of $\eta$ is a regular function on the open set ${T}(L)$. It might or it might not extend to a regular function on the entire $\mathfrak{M}({\Lambda,\mathfrak{t}})$. There are instances where such local merodromies do not extend globally, cf. [CGG+22, Section 11]. Our main result allows us to establish a criterion for such a global extension to exist. It reads as follows:
Theorem 1.1.
@@
Then, for any $\mathscr{D}$-positive relative 1-cycle $\eta$ in $(L,T)$, there exists a Hochschild 0-cycle $H_{\eta}\in\operatorname{HH}_{0}(\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2}))$ whose associated regular function $\operatorname{HO}(H_{\eta})\in\mathrm{H}^{0}\Gamma(\mathfrak{M}({\Lambda,\mathfrak{t}}),\mathcal{O}_{\mathfrak{M}({\Lambda,\mathfrak{t}})})$ coincides with the trace of the microlocal merodromy along $\eta$ when restricted to ${T}(L)\subset\mathfrak{M}({\Lambda,\mathfrak{t}})$.
@@
Then, for any $\mathscr{D}$-positive relative 1-cycle $\eta$ in $(L,T)$, the trace of its microlocal merodromy $A_{\eta}\in\mathrm{H}^{0}\Gamma({T}(L),\mathcal{O}_{{T}(L)})$ extends to a global regular function $A_{\eta}\in\mathrm{H}^{0}\Gamma(\mathfrak{M}({\Lambda,\mathfrak{t}}),\mathcal{O}_{\mathfrak{M}({\Lambda,\mathfrak{t}})})$.
Corollary 1.2 implies the core of the results in [CW24] and [CGG+22], without essentially performing any computations or relying on specific combinatorics. Indeed, the cluster variables defined in [CW24, Section 4] and [CGG+22, Section 5] are particular instances of microlocal merodromies along positive relative cycles, as the relative cycles $\eta$ are the Poincaré duals of the basis given by Lusztig cycles. In addition, [CG24] implies that for these cluster schemes, all elements of the theta canonical basis from [GHKK18] come from such microlocal merodromies. Corollary 1.2 readily establishes their global regularity via a verifiable local condition. Note that such extensions (to the corresponding irreducible component) are unique because $T(L)$ is an open subset of $\mathfrak{M}({\Lambda,\mathfrak{t}})$. We refer to holonomies along positive cycles as positive holonomies: the title of this article thus records the statement of Corollary 1.2.
@@
Note that the microlocal merodromies appearing in [CW24, Section 4] and [CGG+22, Section 5] are a priori only rational functions on the coordinate rings in the ambient space. Key parts of those articles consist of algebraically manipulating such rational expressions, guided by the underlying combinatorics, to prove they are in fact regular functions when restricted to braid varieties. $\Box$
@@
Corollary 1.2 has the important advantage of avoiding the computations mentioned in Remark 1.3. In particular, Theorem 1.1 and Corollary 1.2 apply to arbitrary microlocal ranks and arbitrary Legendrian links (in fact, Legendrian submanifolds in arbitrary dimensions), not just to the moduli space of microlocal rank 1 sheaves on positive braid closures used in [CGG+22, CW24].

### v2

@@ 1. Introduction
Symplectic topology is the study of smooth functions and their first derivatives, broadly understood. Infinitesimally, symplectic geometry studies the differential structure associated to the Lie algebra $\mathfrak{sp}_{2n}$, characterized as the geometry that preserves a non-degenerate skew-symmetric bilinear form. Along with the geometry that preserves volume, corresponding to ${\mathfrak{sl}}_{n}$}, and (symmetric) inner products, corresponding to ${\mathfrak{so}}_{n}$}, these constitute the three infinite families in the classification of complex simple Lie algebras. Locally, symplectic geometry adds the integrability condition that the non-degenerate 2-form is closed, resulting in a unique normal form near any point, up to symmetries preserving the symplectic 2-form. Globally, the intricacies of symplectic 2-forms and their symmetries remain an active subject of study, often referred to as symplectic topology. A salient instance of symplectic manifolds are cotangent bundles $T^{*}Q$ of smooth manifolds $Q$, whereby the smooth topology of $Q$ and its submanifolds is in part reflected by the intersection theory of Lagrangian submanifolds in $T^{*}Q$. Here, Lagrangian submanifolds are defined to be those submanifolds where the restriction of the ambient symplectic 2-form vanishes identically, and are of maximal dimension. In general, such Lagrangian submanifolds generalize the graphs $\mbox{gr}(df)\subset T^{*}Q$ of differentials of smooth functions $f:Q\longrightarrow\mathbb{R}$, in particular allowing for non-graphical behavior and parametric families of such functions.
In modern symplectic topology, Lagrangians and their intersection theory yield useful global invariants of symplectic manifolds. In addition, any closed symplectic manifold admits a symplectic divisor whose complement can be appropriately understood as the cotangent bundle of a possibly singular Lagrangian skeleton. The study of Lagrangian submanifolds is thus of central importance in our understanding of symplectic topology, e.g.  the classification of Lagrangian submanifolds, up to either Hamiltonian isotopies or symplectomorphisms, is a desired goal in the current scientific context. This article examines one of the simplest models: understanding embedded exact Lagrangian surfaces of the local normal form $(\mathbb{R}^{4},\omega_{\text{st}})$, with a fixed boundary condition given by a Legendrian knot. Intuitively, the main goal of the article is to find global regular functions on the derived moduli stack of such Lagrangian surfaces. Such functions allow for a better understanding of these intricate spaces (e.g. sometimes allowing for the computation of their cohomology rings) and yield connections to other areas, including the study of cluster algebras. The main strategy for constructing such global functions is to first construct local regular functions, by using certain parallel transports, and then establish a criterion for these locally regular functions to be globally regular. The title of the manuscript refers to the criterion we present: if such a local function is positive with respect to a Lagrangian skeleton, then it extends to a global regular function.
1.1. Summary
@@
Specifically, our main result is to show that such regular extensions exist if the relative cycle intersects positively an $\mathbb{L}$-compressing system. It applies to arbitrary Legendrian links and Lagrangian fillings (and arbitrary microlocal rank), both generalizing and independently recovering our previous results on global regularity of cluster variables. A key part of our argument is the construction of Hochschild cycles for the aforementioned category that lift these microlocal merodromies. The proof highlights the advantages of working within the dg-categorical framework when studying Legendrian links. Even in known cases, our argument bypasses previously required algebraic computations needed to argue for regularity and, using the derived stack of pseudo-perfect objects, provides a clearer categorical argument for regularity. In brief, apart from its greater level of generality, a strength of the result is that it provides new conceptual and geometric reasons for the global regularity of such functions.
@@ 1.2. Scientific context
Let $\Lambda\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{\text{st}})$ be a Legendrian link in the ideal contact boundary of the cotangent bundle of $\mathbb{R}^{2}$, e.g. any Legendrian link in a contact Darboux chart. The study of Lagrangian fillings of $\Lambda$ is a pillar of low-dimensional contact topology, see e.g. [BST15, EEN+13, EHK16, EP96, EN02, GKS12, LEV16, NRS+20, POL91, STZ17]. Recently, the results from [CG22, CG24] and the trilogy [CGG+25, CW24, CZ22] provided a better understanding of the moduli space $\mathfrak{M}(\Lambda)$ of Lagrangian fillings and its geometric structures. Note that, in addition to new results in contact topology, studying these moduli spaces $\mathfrak{M}(\Lambda)$ from this contact topological perspective has lead to a number of applications to cluster algebras, see e.g. [CGG+24, CGG+25, CLS+23, CW24]. Now, a key technique to understand such moduli spaces $\mathfrak{M}(\Lambda)$ is the construction of regular functions on them. The main result of this article is a general construction of such regular functions coming from a microlocal version of parallel transport.
In a nutshell, from [CW24, Section 4], certain relative 1-cycles $\eta\in H_{1}(L,\Lambda)$ in a Lagrangian filling $L$ (related to $\mathbb{L}$-compressing systems) allow us to define regular functions $A_{\eta}:T(L)\longrightarrow\mathbb{C}$ on an open subset $T(L)\subset\mathfrak{M}(\Lambda)$. If $\Lambda$ is particular enough and certain combinatorics are present, e.g. weaves in the case of [CGG+25], grid plabic graphs in the case of [CW24] and 3D plabic graphs in the case of [GLS+22], one can sometimes argue that $A_{\eta}$ extends to a regular function on the entirety of $\mathfrak{M}(\Lambda)$. This has significant consequences, including the construction of cluster structures [CGG+25, CW24], holomorphic symplectic structures [CGG+24], the existence of Lagrangian fillings in each cluster seed [CG24] and geometric realizations of Donaldson-Thomas transformations [CLS+23, CW24].
@@ 1.3. Main result
Let $\Lambda\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{st})$ be a Legendrian link and $\mathfrak{t}\subset\Lambda$ a set of basepoints, with at least one basepoint per component of $\Lambda$. We say $\Lambda$ is $\mathfrak{t}$-pointed if such a choice $\mathfrak{t}$ of basepoints has been made. In Section 2 we introduce a refinement $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ of the dg-category $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ of compactly supported sheaves in $\mathbb{R}^{2}$ with singular support on $\Lambda$. The objects of $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ are categorically compact sheaves in $\operatorname{Sh}_{\Lambda}(\mathbb{R}^{2})_{0}$ together with trivializations of their microstalks at the basepoints $\mathfrak{t}$. By construction, $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ is a smooth dg-category and we denote by $\mathfrak{M}({\Lambda,\mathfrak{t}})$ its moduli of pseudo-perfect objects, cf. [TV07, Section 3.1]. The derived stack $\mathfrak{M}({\Lambda,\mathfrak{t}})$ is the model we use for the moduli space of Lagrangian fillings of $(\Lambda,\mathfrak{t})$: intuitively, its classical closed points parametrize Lagrangian fillings of $\Lambda$ endowed with local systems, trivialized at some boundary points. We will use the data of the microstalk trivializations in $\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})_{0}$ to obtain regular functions on $\mathfrak{M}({\Lambda,\mathfrak{t}})$, by using the parallel transport of the local systems along relative cycles. For precise details, see e.g. [CL23, CW24] and [TV07]. Now, the construction of [TV07] is such that there exists a map
@@
Let $T\subset\Lambda\setminus\mathfrak{t}$ be another set of basepoints such that there exists exactly one basepoint of $\mathfrak{t}$ in each interval of $\Lambda\setminus T$. We say $\Lambda$ is $(\mathfrak{t},T)$-pointed if such choices of basepoints $\mathfrak{t}$ and $T$ have been made. Consider an embedded exact Lagrangian filling $L\subset(T^{*}\mathbb{R}^{2},\lambda_{\text{st}})$ of $\Lambda$ endowed with an $\mathbb{L}$-compressing system $\mathscr{D}=\{\gamma_{1},\ldots,\gamma_{b_{1}(L)}\}$, as defined in [CG24, Section 4]. Succinctly, a (complete) $\mathbb{L}$-compressing system is a set of disjoint Lagrangian disks $\{D_{1},\ldots,D_{b_{1}(L)}\}$, each properly embedded in the complement $T^{*}\mathbb{R}^{2}\setminus L$ of $L$ with immersed boundaries $\gamma_{i}:=\partial\overline{D_{i}}\subset L$ and such that the union $\mathbb{L}:=L\cup(D_{1}\cup\ldots D_{b_{1}(L)})\subset T^{*}\mathbb{R}^{2}$ is a relative Lagrangian skeleton for $(T^{*}\mathbb{R}^{2},\Lambda)$. See also [CW24, Section 1]. [^1] The assumption that $\mathbb{L}:=L\cup(D_{1}\cup\ldots D_{b_{1}(L)})\subset T^{*}\mathbb{R}^{2}$ is a relative Lagrangian skeleton for $(T^{*}\mathbb{R}^{2},\Lambda)$ is not explicitly stated in [CW24, Section 1], but all known examples of $\mathbb{L}$-compressing systems, in the sense of [CW24], satisfy this assumption. See for example [CAS22] and [BRW25, Proposition 5.2] for a proof. By definition, a relative cycle $\eta:[0,1]\longrightarrow(L,\mathfrak{t})$ is said to be $\mathscr{D}$-positive if $\langle\eta,\gamma_{i}\rangle\geq 0$ for all $i\in[1,b_{1}(L)]$.
Let ${T}(L)\subset\mathfrak{M}({\Lambda,\mathfrak{t}})$ be the open set ${T}(L)\cong\mathrm{Loc}(L,T)$ given by the embedded exact Lagrangian filling $L$, see e.g. [CL23, Appendix B] or [CW24, Section 2] and references therein. By construction, the microlocal monodromy $A_{\eta}\in\Gamma({T}(L),\mathcal{O}_{{T}(L)})$ of $\eta$ is a regular function on the open set T(L){T}(L). It might or it might not extend to a regular function on the entire $\mathfrak{M}({\Lambda,\mathfrak{t}})$. There are instances where such local merodromies do not extend globally, cf. [CGG+25, Section 11]. Our main result allows us to establish a criterion for such a global extension to exist. It reads as follows:
Theorem 1.1 (“Main extension result”).
@@
Then, for any $\mathscr{D}$-positive relative 1-cycle $\eta$ in the relative pair $(L,T)$, there exists a Hochschild 0-cycle $H_{\eta}\in\operatorname{HH}_{0}(\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2}))$ whose associated regular function $\operatorname{HO}(H_{\eta})\in\mathrm{H}^{0}\Gamma(\mathfrak{M}({\Lambda,\mathfrak{t}}),\mathcal{O}_{\mathfrak{M}({\Lambda,\mathfrak{t}})})$ coincides with the trace of the microlocal merodromy along $\eta$ when restricted to ${T}(L)\subset\mathfrak{M}({\Lambda,\mathfrak{t}})$.
@@
Then, for any $\mathscr{D}$-positive relative 1-cycle $\eta$ in the relative pair $(L,T)$, the trace of its microlocal merodromy $A_{\eta}\in\mathrm{H}^{0}\Gamma({T}(L),\mathcal{O}_{{T}(L)})$ extends to a global regular function $A_{\eta}\in\mathrm{H}^{0}\Gamma(\mathfrak{M}({\Lambda,\mathfrak{t}}),\mathcal{O}_{\mathfrak{M}({\Lambda,\mathfrak{t}})})$.
Corollary 1.2 implies the core of the results in [CW24] and [CGG+25], without essentially performing any computations or relying on combinatorics. Indeed, the cluster variables defined in [CW24, Section 4] and [CGG+25, Section 5] are particular instances of microlocal merodromies along positive relative cycles, as the relative cycles $\eta$ are the Poincaré duals of the basis given by Lusztig cycles. In addition, [CG24] implies that for these cluster schemes, all elements of the theta canonical basis from [GHK+18] come from such microlocal merodromies. Corollary 1.2 readily establishes their global regularity via a verifiable local condition. Note that such extensions (to the corresponding irreducible component) are unique because $T(L)$ is an open subset of $\mathfrak{M}({\Lambda,\mathfrak{t}})$. We refer to holonomies along positive cycles as positive holonomies: the title of this article thus records the statement of Corollary 1.2.
@@
The microlocal merodromies appearing in [CW24, Section 4] and [CGG+25, Section 5] are a priori only rational functions on the coordinate rings in the ambient space. Key parts of those articles consist of algebraically manipulating such rational expressions, guided by the underlying combinatorics, to prove they are in fact regular functions when restricted to braid varieties. $\Box$
@@
Corollary 1.2 has the important advantage of avoiding the computations mentioned in Remark 1.3. In particular, Theorem 1.1 and Corollary 1.2 apply to arbitrary microlocal ranks and arbitrary Legendrian links in contact boundaries, not just to the moduli space of microlocal rank 1 sheaves on positive braid closures used in [CGG+25, CW24], cf. Section 4.6.
1.4. Outline of the argument
The proof of 1.1 is rather conceptual and it can be helpful to emphasize the three key ingredients that feature in the proof. We thank the referee for suggesting that we add the following list in this introduction. The main ingredients can be summarized as follows:
  1. (1)
    The moduli of pseudoperfect objects $\mathfrak{M}_{\mathcal{C}}$ of a smooth dg-category $\mathcal{C}$, as introduced and studied by B. Toën and M. Vaquié in [TV07]. By [TV11, Corollary 4.2], the dg-category $\mathcal{C}$ and its derived stack $\mathfrak{M}_{\mathcal{C}}$ are equipped with a map $\operatorname{HO}:\operatorname{HH}_{*}(\mathcal{C})\longrightarrow\Gamma(\mathfrak{M}_{\mathcal{C}},\mathcal{O}_{\mathfrak{M}_{\mathcal{C}}})$ from the Hochschild chains of $\mathcal{C}$ to the derived global functions on $\mathfrak{M}_{\mathcal{C}}$. Therefore, to construct functions on $\mathfrak{M}_{\mathcal{C}}$ it suffices to construct Hochschild chains on $\mathcal{C}$. In the proof of 1.1, this general framework is applied to the dg-category $\mathcal{C}:=\operatorname{Sh}^{c}_{\Lambda,\mathfrak{t}}(\mathbb{R}^{2})$ that we introduce in Section 2.1, and the necessary details on $\mathfrak{M}_{\mathcal{C}}$ and the $\operatorname{HO}$ map are provided in Section 2.2 and Section 4.1, respectively.
  2. (2)
    The local model of the Lagrangian skeleton near a positive path, as depicted in Fig. 2. Given a generic relative chain in the union of a Lagrangian filling and an $\mathbb{L}$-compressing system, such as the relative chain $\eta$ drawn in red in Fig. 1, its neighborhood retracts to a horizontal line (or circle) with a sequence of vertical half-rays attached to it, some pointing upwards and some downwards. For a positive path, these half-rays all point upwards, as in Fig. 2. This local model is studied in detail in Section 4.2, where Hochschild chains for its associated dg-category are computed, and in Section 4.3, where the corresponding $\operatorname{HO}$ map is described in 4.5.
  3. (3)
    The corestriction functors and their compatibility with Hochschild chains, as studied in Sections 3.2, 3.3 and 4.4. The existence of corestriction functors for the Kashiwara-Schapira stack is established in Section 3.2, and in the pointed setting in Section 3.3. These corestriction functors, which are left adjoints to restriction functors, allow for a local-to-global transition from the local computation in the local model from (2) above, near a relative chain, to a global computation on the Lagrangian skeleton. Section 4.4 proves that such functors are compatible with Hochschild chains and the $\operatorname{HO}$ maps, thus allowing for the local computation of the $\operatorname{HO}$ map to extend globally.
Having established the needed results for each of these three main ingredients in Sections 2, 3 and 4, 1.1 is proven in Section 4.5.